I have 10% of homemade salsa that is 25% jalapenos. I buy some store salsa that is 4% and mix the two salsa together. how much store salsa do i need in order to have a mixture that is 10% jalapenos?
Answer by stanbon(75887)

I have 10% of homemade salsa that is 25% jalapenos. I buy some store salsa that is 4% and mix the two salsa together. how much store salsa do i need in order to have a mixture that is 10% jalapenos?
--------------------------
Equation:
jalapeno + jalapeno = jalapeno
0.25*10 + 0.04x = 0.10(10+x)
---
Multiply thru by 100 to get:
25*10 + 4x = 10*10+10x
6x = 15*10
x = 25 (amount of 4% salsa needed in the mixture)
=======================================================
Cheers,
Stan H.